OpenStudy (jason_danford):

What is the velocity of an Earth satellite that is in a circular orbit with a radius of 7.5 × 107 meters measured from the center of the Earth?

OpenStudy (irishboy123):

it's mass independent. balancing radial force and inertia: \(F = \dfrac{GMm}{r^2} = \dfrac{mv^2}{r}\) \( \dfrac{GM}{r} = v^2\) and you can google G and M